In a triangle ABC prove that cos A / cos C = c-acosB/ a-cCos B

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
By the projection formula:

c = bcosA + acosB  and a = bcosC + ccosB

Therefore, c - a cosB = b cosA and a - c cosB = b cosC

Therefore, (c - a cosB)/(a - c cosB) = cosA/cosC
 
Hence proved
 
 
Regards
Arun (askIITians forum expert)